# Pairing in Fission: Studies of Static Approach and Collective Inertias

### v1

@@ III RESULTS
Figure 5: The logarithms of spontaneous fission half-lives of 250Cm, 260Fm and 262Rf isotopes obtained within different values of the paring strength and compared with the experimental data taken from Kondev et al. (2021) depicted as horizontal lines.
@@
Our first goal is to understand the impact of the pairing gap, particle number fluctuations, and the imposed pairing strength on the spontaneous fission static description. In order to clarify the picture and to minimize the influence of the secondary effects by involving the parity breaking, we decided to focus first on the case of symmetric fission. We have chosen for our analysis the 262Rf isotope for which spontaneous fission half-life, as well as the fission fragment mass yields, are measured and available for comparison.
@@ III.1 The pairing gap
As has already been emphasized in the Introduction, the first studies within a simple, realistic model regarding the consideration of pairing in the description of fission showed that an increase in the pairing gap parameter leads to an increase in the penetrability of the fission barrier. This effect is strictly related to the fact that the collective inertias decrease as the inverse square of the pairing gap Moretto and Babinet (1974). Panel (a) of Fig. 1 depicts the potential energy surface (PES) in the ($Q_{20},\Delta$) plane of 262Rf. The solid white line represents the least-energy path in this space of the collective coordinates. To facilitate an understanding of the influence of $\Delta$ on the potential energy of the nucleus, we have also created a one-dimensional plot of the potential energy (panel b) and pairing energy (panel c) as a function of nuclear elongation. The black line corresponds to the path of least energy, while the other colors represent changes in energy as a function of $Q_{20}$ for a fixed value of the pairing gap greater than what was found in one-dimensional calculations with the constraint only on the quadrupole moment. It is evident that as $\Delta$ increases, the energy of the system increases as well. As expected, this is accompanied by an increase in the absolute value of pairing energy, as shown in panel (b), and a significant decrease in the values of collective inertias, as shown in panel (d). Apart from that, one should notice that the fission barrier becomes gradually broader for higher $\Delta$ values.
@@ III.2 The particles number fluctuation
The particle number fluctuations $\Delta\hat{N}^{2}$, which is a quantity proportional to the pairing gap, has been successfully used for estimations of the spontaneous fission half-lives of actinides Sadhukhan et al. (2014); Giuliani et al. (2014); Zhao et al. (2016); Rodríguez-Guzmán et al. (2023). Most of the studies found that the coupling of $\Delta\hat{N}^{2}$ with shape degrees of freedom greatly improves the agreement between theoretical and experimental results. Now, we have decided to explore in detail how the particle number fluctuations as a collective degree of freedom impacts the fission barriers. The panel (a) of Fig. 2 presents the PES of 262Rf in the ($Q_{20},\Delta\hat{N}^{2}$) plane. One can notice that the topography of the presented surface looks very similar to Fig. 1. Also, the behavior of the analyzed quantities presented in panels (c) and (d) is similar to the one observed when the pairing gap was employed as a collective degree of freedom. This is a direct consequence of the almost linear connection between $\Delta$ and $\Delta\hat{N}^{2}$ variables. As $\Delta\hat{N}^{2}$ increases, the potential energy increases and the collective inertia decreases, both following a quadratic trend. As expected, in both cases, whether the pairing gap or the particle number fluctuations increases, it is accompanied by the broadening of the fission barrier As the energy and collective inertia profiles obtained with $\Delta$ and $\Delta\hat{N}^{2}$ show a very similar pattern, it is possible to conclude that these two quantities may be treated interchangeably.
@@ III.3 The quenching factor
By considering different values for the quenching factor, one can scale the HFB proton and neutron pairing fields, which in practice leads to an increased or decreased pairing strength with respect to the value found in the process of fitting the parameters of the effective interaction. It has been shown that by increasing the quenching factor by 10-20%, one can obtain a better agreement with the data for the typical fission observable, like spontaneous fission half-lives of even-even Giuliani and Robledo (2013); Rodríguez-Guzmán and Robledo (2014a, b) and odd-mass Guzman and Robledo (2017) systems as well as fission fragment mass distributions Guan et al. (2023); Zhao et al. (2021). Fig. 3 depicts the least-energy path of 262Rf as a function of elongation for various values of the quenching factor. The presented PES is plotted in relation to the path obtained with $\textup{QF}=1$. Since the quenching factor is one of the three microscopic quantities related to pairing, one would expect that changes in the value of this parameter would lead to similar effects on energy and collective inertias as the previously analyzed $\Delta\hat{N}^{2}$ and $\Delta$. However, the results of the calculations conducted for the variable QF demonstrate that we are dealing with a different effect here, as we are changing the wave function by changing the underlying interaction. It should be noted that as the quenching factor increases, the potential energy of the nucleus decreases as the attractive pairing interaction becomes stronger. This is the opposite effect compared to what was observed earlier for increasing $\Delta\hat{N}^{2}$ and $\Delta$, which caused an increase in the potential energy. It is also worth noting that while increasing the particle number fluctuations, and the pairing gap results in higher and broader fission barrier, in the case of an increase in QF, the barrier becomes lower while its width does not change significantly. Furthermore, as shown in Fig. 3, further increasing the value of QF results in a monotonic decrease in the potential energy of the nucleus and a flattening of the fission barrier.
One can conclude that the quenching factor can be used as an artefact to renormalise locally the pairing strength in order to better reproduce experimental data. It can be also used for testing fission with enhanced pairing along the fission path. However, this method does not allow to dynamically adjust pairing strength for the search of the least action path.
@@ III.4 The quenching factor and spontaneous fission half-lives evaluation
The usual approach to describe the spontaneous fission process is to follow the least-energy path. This is a reasonable approximation that produces qualitative and often quantitative results in good agreement with experimental data. By changing the quenching factor, one can expect a slightly different topography of the PES, as discussed in Sec. III.3. The upper panel of the Fig. 4 presents the fission barrier profiles corrected by the zero-point energy (ZPE) for 262Rf (left) and 260Fm (right). We notice a reduction of the fission barrier by a few MeV when the pairing strength is increased by 10%. The collective inertias $B_{22}$ are plotted in the middle panels of Fig. 4. Their values decrease together with increasing the quenching factor and become smoother. With the stronger pairing, the difference in the $B_{22}$ values becomes less apparent. The bottom panel of Fig. 4 presents the pairing energies of protons and neutrons. Obviously, the increase of the quenching factor leads to an increase of the absolute value of pairing energy. Fig. 5 presents the logarithms of spontaneous fission half-lives resulting from the static studies for various values of QF. As expected from the above analysis, the decrease of the fission barriers and the collective inertias leads to an increase of the probability of tunneling the fission barrier, leading to shorter half-lives as the pairing strength grows.
@@
Let us now take a look at the actinides in order to have some insight into the pairing properties at the asymmetric fission path. For our studies, we have chosen the plutonium and uranium isotopic chains, where the asymmetric fission mode exists. At the panel (a) of Fig. 6, the least energy paths of two representative isotopes, namely 238Pu and 236U, are plotted for various values of QF. Here, the first hump of the barrier is reflection symmetric, whereas the second one is octupolly deformed. One can see that the increased QF lowers the first symmetric fission barrier by approximately 0.4 MeV per 0.05 change in the QF value. The second minimum and the second asymmetric fission barrier show a slightly smaller influence of pairing strength and the diminishing barrier height between the considered QF values does not exceed 0.2 MeV. This trend is inversed behind the second barrier, at around $Q_{20}$ = 80 b. Beyond this point, the barrier becomes higher for a higher value of the quenching factor. The collective inertias related to the discussed fission paths are presented in the lower panel of Fig. 6. Similarly, as for the previously discussed symmetric fission, $B_{22}$ is reduced and smoothened out for higher QF values. It is worth to mention that in these calculations, despite the second barrier in this region of actinides leads through the octupole shapes, the components of the inertia tensor associated with this degree of freedom have been neglected.
@@
The calculated spontaneous fission half-lives are plotted in Fig 7. Looking at the results, one can immediately notice that the optimal value of the pairing strength for reproducing experimental half-lives seems to be between 105% and 110% for the D1S parametrization. The effect of the increased tunneling probability when the quenching factor increases, and consequently the reduction of the spontaneous fission half-lives, mostly comes from the action integral of the first hump of the fission barrier. This is because of the strong reduction of collective inertias with higher QF, accomplished by diminishing the potential energy. The influence of the asymmetric part of the barrier on the half-lives with increasing QF is less important. First, despite this part of the barrier being relatively broad, the collective inertia is much smaller than at the first barrier. Second, the reduction of the potential energy around the peak is compensated by an increase of the energy in the tail of the barrier. It means that the action integral is insensitive to the energy changes introduced by the QF in the asymmetric part of the least-energy path. This suggests that QF has a stronger impact when the parity is not broken. Thus, the overall trend of the shorter half-lives for higher QF in the actinides mainly comes from the response of the symmetric barrier on the pairing strength value.
@@ III.5 Collective inertias with constraints on $\Delta$ and $\Delta\hat{N}^{2}$
Figure 8: The effective inertia $B(s)$ along the quadrupole moment $s=Q_{20}$ for several values of $\Delta$ (solid line) and the corresponding $\Delta\hat{N}^{2}$ (dashed line) for 262Rf - see the text for detailed explanation.
Further investigation of the role of microscopic pairing in fission should involve a dynamic approach requiring the calculation of the collective inertia for the pairing modes. Among the available quantities associated with the pairing correlations discussed in this paper, the quenching factor can not be considered a dynamical degree of freedom as it involves the change of the interaction itself, and it is, therefore hardly consistent with the traditional frameworks to define collective inertias. On the other hand, the $\Delta$ and $\Delta\hat{N}^{2}$ collective variables can be used for that purpose, and it is interesting to compare the inertias associated with both of them. As already mentioned in Section II, $\Delta\hat{N}^{2}$ is a two-body operator, which means that the standard cranking approximation cannot be applied straightforwardly. Following the traditional procedure, we decided to check the collective inertia associated with this degree of freedom when one uses in the formulas for the inertia the quasiparticle one body part of $\Delta\hat{N}^{2}$. In order to check the validity of this assumption, we compared the resulting $\Delta\hat{N}^{2}$ inertias with those associated to $\Delta$. The effective inertia is defined in terms of the collective inertia tensor $B_{ij}$ along the fission path $s=Q_{20}$:
@@
where $B_{ij}$ is the inverse of the collective mass tensor calculated within the perturbative cranking approximation Baran et al. (2011), and $(q_{i},q_{j})$ stand for the collective variable pair $(Q_{20},\Delta)$ and $(Q_{20},\Delta\hat{N}^{2})$. We have considered several fixed values of $\Delta$ and related to their values of the $\Delta\hat{N}^{2}$ for which the collective inertias along the quadrupole moment have been computed. The results for 262Rf are shown in Fig. 8. The solid lines represent the inertias associated with $\Delta$ while the dashed lines correspond to the ones obtained when $\Delta\hat{N}^{2}$ is treated as collective variable. As one can see, the effective inertias associated with $\Delta\hat{N}^{2}$ agree very well with the $\Delta$-dependent ones. It is worth emphasizing that the behavior of both inertias becomes more consistent as the value of $\Delta$ and $\Delta\hat{N}^{2}$ increases. This is particularly important in dynamic calculations, as one should expect that the path of the least-action path will primarily pass through higher values of these constraints. These results suggest that, in first approximation, one can apply the one-body expression for the inertias in the perturbative cranking approximation employing the quasiparticle one body part of $\Delta\hat{N}^{2}$ as a dynamical variable. Developments in this direction are currently in progress.
@@ IV CONCLUSIONS
In this work, we analyzed the different methods traditionally employed in studies exploring the impact of pairing correlations in the static description of spontaneous fission. We compared the potential energy surfaces, fission barrier, collective inertias, and pairing energies obtained when $\Delta$ and $\Delta\hat{N}^{2}$ are employed as collective degrees of freedom. The results have been confronted with those obtained when a quenching factor QF is applied to the HFB pairing field.
@@
The strength of pairing interaction along the fission path from ground state to scission can be controlled in various ways. One may increase the absolute value of the pairing energy by raising the pairing gap ($\Delta$), the particle number fluctuations ($\Delta\hat{N}^{2}$), or the pairing strength.
@@
Conversely, the increase of the pairing strength by scaling the quenching factor leads to a decrease of the potential energy. Lowering the fission barrier and the collective inertias simultaneously leads to a monotonic reduction of the fission half-lives for stronger pairing interactions. The effect of lowering the barrier with an increase of quenching factor is less significant in the region of the asymmetric fission channel. Thus, enhancing the pairing strength has a different impact on fission properties compared to the two previous techniques. This is because the nuclear interactions are modified. Changes in $\Delta$ and $\Delta\hat{N}^{2}$ touches only the single-particle level properties around the Fermi level, but the nuclear force remains unchanged.
The effective inertias associated with the $\Delta\hat{N}^{2}$ computed within the “one-body assumption" are in a very good agreement with the ones obtained within a standard formula applicable for one-body operators, giving a solid ground to pursue the comparative dynamical studies.

### v2

@@ III RESULTS
Figure 5: The logarithm of spontaneous fission half-live of 250Cm, 260Fm and 262Rf isotopes obtained within different values of the paring strength and compared with the experimental data taken from Kondev et al. (2021) depicted as horizontal lines.
@@
Our first goal is to understand the impact of the pairing gap, particle number fluctuations, and the imposed pairing strength on the spontaneous fission static description. In order to clarify the picture and to minimize the influence of the secondary effects by involving parity breaking, we decided to focus first on the case of symmetric fission. We have chosen for our analysis the 262Rf isotope for which spontaneous fission half-life, as well as the fission fragment mass yields, have been measured and are available for comparison.
@@ III.1 The pairing gap
As has already been emphasized in the Introduction, the first studies within a simple, realistic model studying the influence of pairing in the description of fission showed that an increase in the pairing gap parameter leads to an increase in the penetrability of the fission barrier. This effect is strictly related to the fact that the collective inertias decrease as the inverse square of the pairing gap Moretto and Babinet (1974).
Panel (a) of Fig. 1 depicts the potential energy surface (PES) in the ($Q_{20},\Delta$) plane of 262Rf. The solid white line represents the least-energy path in this space of collective coordinates. To facilitate the understanding of the influence of $\Delta$ on the potential energy of the nucleus, we have also created one-dimensional plots of the potential energy (panel b) and pairing energy (panel c) as a function of nuclear elongation. The black line corresponds to the values along the least-energy path, while the other colors are associated with calculations with a constrained value of $\Delta$. It is evident that as $\Delta$ increases, the energy of the system increases as well. As expected, this is accompanied by an increase in the absolute value of pairing energy, as shown in panel (b), and a significant decrease in the values of collective inertias, as shown in panel (d). The decrease in the inertia follows quite closely the inverse of the square of the pairing gap law as it can be checked that the inertias for $\Delta=2.4$ are a factor 1.8 larger than the ones of $\Delta=3.2$ over the whole range of quadrupole moments. Apart from that, one should also notice that the fission barrier becomes gradually broader for higher $\Delta$ values.
@@ III.2 The particles number fluctuation
The particle number fluctuations $\Delta\hat{N}^{2}$, which is a quantity proportional to the pairing gap, have been successfully used to estimate the spontaneous fission half-lives of actinides Sadhukhan et al. (2014); Giuliani et al. (2014); Zhao et al. (2016); Rodríguez-Guzmán et al. (2023). Most of the studies found that the coupling of $\Delta\hat{N}^{2}$ with shape degrees of freedom greatly improves the agreement between theoretical and experimental results. Now, we have decided to explore in detail how the particle number fluctuations as a collective degree of freedom impact the fission barriers. The panel (a) of Fig. 2 presents the PES of 262Rf in the ($Q_{20},\Delta\hat{N}^{2}$) plane. One can notice that the topography of the presented surface looks very similar to Fig. 1. Also, the behavior of the analyzed quantities presented in panels (c) and (d) is similar to the one observed when the pairing gap was employed as a collective degree of freedom. This is a direct consequence of the almost linear connection between $\Delta$ and $\Delta\hat{N}^{2}$ variables (see Fig. 9 in the Appendix). As $\Delta\hat{N}^{2}$ increases, the potential energy increases and the collective inertia decreases, both following a quadratic trend as can be observed in Fig 1 in Ref Giuliani et al. (2014). It is also observed that an increase in the particle number fluctuations is accompanied by the broadening of the fission barrier. As the energy and collective inertia profiles obtained with $\Delta$ and $\Delta\hat{N}^{2}$ show a very similar pattern, it is possible to conclude that these two quantities may be treated interchangeably as collective variables. To test this conclusion, we have performed calculations of the overlap between the HFB states obtained with the constraint on $\Delta\hat{N}^{2}$ and that obtained with constraint on $\Delta$ to the value obtained with the $\Delta\hat{N}^{2}$ constraint. The values of the overlaps are always very close to one.
@@ III.3 The quenching factor
By considering different values for the quenching factor, one can scale the HFB proton and neutron pairing fields, which in practice leads to an increase or decrease of the pairing strength with respect to the nominal value found in the process of fitting the parameters of the effective interaction. It has been shown that by increasing the quenching factor by 10-20%, one can obtain a better agreement with the data for the typical fission observables, like spontaneous fission half-lives of even-even Giuliani and Robledo (2013); Rodríguez-Guzmán and Robledo (2014a, b) and odd-mass Guzman and Robledo (2017) systems as well as fission fragment mass distributions Guan et al. (2023); Zhao et al. (2021). Fig. 3 depicts the least-energy path of 262Rf as a function of elongation for various values of the quenching factor. The presented PES is plotted in relation to the path obtained with $\textup{QF}=1$. Since the quenching factor is one of the three variables related to pairing, one would expect that changes in the value of this parameter would lead to similar effects on energy and collective inertias as in the previously analysed $\Delta\hat{N}^{2}$ and $\Delta$ cases. However, the results of the calculations conducted for the variable QF demonstrate that we are dealing with a different effect here, as we are changing the wave function by changing the underlying interaction. It should be noted that as the quenching factor increases, the potential energy of the nucleus decreases as the attractive pairing interaction becomes stronger. This is the opposite effect compared to what was observed earlier for increasing $\Delta\hat{N}^{2}$ and $\Delta$, which caused an increase in the potential energy. It is also worth noting that while increasing the particle number fluctuations, and the pairing gap results in higher and broader fission barrier, in the case of an increase in QF, the barrier becomes lower while its width does not change significantly. Furthermore, as shown in Fig. 3, further increasing the value of QF results in a monotonic decrease in the potential energy of the nucleus and a flattening of the fission barrier.
One can conclude that the quenching factor can be used as an artifact to renormalise locally the pairing strength in order to better reproduce experimental data. It can also be used for exploring fission properties with enhanced pairing along the fission path. However, this method does not allow for dynamical adjustment of pairing strength while searching for the least-action path.
@@ III.4 The quenching factor and spontaneous fission half-lives evaluation
The usual approach to describe the spontaneous fission process is to follow the least-energy path. This is a reasonable approximation that produces qualitative and often quantitative results in good agreement with experimental data. By changing the quenching factor, one can expect a slightly different topography of the PES, as discussed in Sec. III.3. The upper panel of Fig. 4 presents the fission barrier profiles corrected by the zero-point energy (ZPE) for 262Rf (left) and 260Fm (right). We notice a reduction of the fission barrier by a few MeV when the pairing strength is increased by 10%. The collective inertias $B_{22}$ are plotted in the middle panels of Fig. 4. Their values decrease together with increasing the quenching factor and become smoother. With the stronger pairing, the difference in the $B_{22}$ values becomes less apparent. The bottom panel of Fig. 4 presents the pairing energies of protons and neutrons. Obviously, the increase of the quenching factor leads to an increase in the absolute value of pairing energy. Fig. 5 presents the logarithms of spontaneous fission half-lives resulting from the static studies for various values of QF. As expected from the above analysis, the decrease of the fission barriers and collective inertias lead to an increase in the probability of tunneling the fission barrier, leading to shorter half-lives as the pairing strength grows.
@@
Let us now take a look at the actinides in order to have some insight into the pairing properties at the asymmetric fission path. For our studies, we have chosen the plutonium and uranium isotopic chains, where the asymmetric fission mode exists. In the panel (a) of Fig. 6, the least-energy paths of two representative isotopes, namely 238Pu and 236U, are plotted for various values of QF. Here, the first hump of the barrier is reflection symmetric, whereas the second one is octupolly deformed. One can see that the increased QF lowers the first symmetric fission barrier by approximately 0.4 MeV per 0.05 change in the QF value. The second minimum and the second asymmetric fission barrier show a slightly smaller influence of pairing strength and the diminishing barrier height between the considered QF values does not exceed 0.2 MeV. This trend is inverted beyond the second barrier, at around $Q_{20}$ = 90 b and energy becomes higher for a higher value of the quenching factor. The collective inertias related to the discussed fission paths are presented in the lower panel of Fig. 6. Similarly, as for the previously discussed symmetric fission, $B_{22}$ is reduced and smoothened out for higher QF values. It is worth mentioning that in these calculations, despite the second barrier in this region of actinides leads through the octupole shapes, the components of the inertia tensor associated with this degree of freedom have been neglected.
@@
The calculated spontaneous fission half-lives are plotted in Fig 7. Looking at the results, one can immediately notice that the optimal value of the pairing strength for reproducing experimental half-lives corresponds to a QF between 105% and 110% for the D1S parametrization. The effect of the increased tunneling probability when the quenching factor increases, and consequently the reduction of the spontaneous fission half-lives, mostly comes from the action integral of the first hump of the fission barrier. This is because of the strong reduction of collective inertias with higher QF, accomplished by diminishing the potential energy. The influence of the asymmetric part of the barrier on the half-lives with increasing QF is less important. First, despite this part of the barrier being relatively broad, the collective inertia is much smaller than at the first barrier. Second, the reduction of the potential energy around the peak is compensated by an increase of the energy in the tail of the barrier. It means that the action integral is insensitive to the energy changes introduced by the QF in the asymmetric part of the least-energy path. This suggests that QF has a stronger impact when parity is not broken. Thus, the overall trend of the shorter half-lives for higher QF in the actinides mainly comes from the response of the symmetric barrier on the pairing strength value.
@@ III.5 Collective inertias with constraints on $\Delta$ and $\Delta\hat{N}^{2}$
Figure 8: The effective inertia $B(s)$ along the quadrupole moment $s=Q_{20}$ for several values of $\Delta$ (solid line) and the corresponding $\Delta\hat{N}^{2}$ (dashed line) for 262Rf - see the text for the detailed explanation.
Further investigation of the role of microscopic pairing in fission should involve a dynamic approach requiring the calculation of the collective inertia for the pairing modes. Among the available quantities associated with the pairing correlations discussed in this paper, the quenching factor can not be considered a dynamical degree of freedom as it involves the change of the interaction itself, and it is, therefore hardly consistent with the traditional frameworks to define collective inertias. On the other hand, the $\Delta$ and $\Delta\hat{N}^{2}$ collective variables can be used for that purpose, and it is interesting to compare the inertias associated with both of them. As already mentioned in Section II, $\Delta\hat{N}^{2}$ is a two-body operator, which means that the standard cranking approximation, defined for one body operators, cannot be applied straightforwardly. Following the traditional procedure, we decided to check the collective inertia associated with this degree of freedom when one uses in the formulas for the inertia of the quasiparticle one body part of $\Delta\hat{N}^{2}$. In order to check the validity of this assumption, we compared the resulting $\Delta\hat{N}^{2}$ inertias with those associated to $\Delta$. The effective inertia is defined in terms of the collective inertia tensor $B_{ij}$ along the fission path $s=Q_{20}$:
@@
where $B_{ij}$ is the inverse of the collective mass tensor calculated within the perturbative cranking approximation Baran et al. (2011), and $(q_{i},q_{j})$ stand for the collective variable pair $(Q_{20},\Delta)$ and $(Q_{20},\Delta\hat{N}^{2})$. We have considered several fixed values of $\Delta$ and related to their values of the $\Delta\hat{N}^{2}$ for which the collective inertias along the quadrupole moment have been computed. The results for 262Rf are shown in Fig. 8. The solid lines represent the inertias associated with $\Delta$ while the dashed lines correspond to the ones obtained when $\Delta\hat{N}^{2}$ is treated as a collective variable. As one can see, the effective inertias associated with $\Delta\hat{N}^{2}$ agree very well with the $\Delta$-dependent ones. It is worth emphasizing that the behavior of both inertias becomes more consistent as the value of $\Delta$ and $\Delta\hat{N}^{2}$ increases. This is particularly important in dynamic calculations, as one should expect that the path of the least-action path will primarily pass through higher values of these constraints. These results suggest that, in a first approximation, one can apply the one-body expression for the inertias in the perturbative cranking approximation employing the quasiparticle one body part of $\Delta\hat{N}^{2}$ as a dynamical variable. Developments in this direction are currently in progress.
@@ IV CONCLUSIONS
In this work, we analyzed the different methods traditionally employed in studies exploring the impact of pairing correlations in the static description of spontaneous fission. We compared the potential energy surfaces, fission barrier, collective inertias, and pairing energies obtained when $\Delta$ and $\Delta\hat{N}^{2}$ are used as collective degrees of freedom. The results have been confronted with those obtained when a quenching factor QF is applied to the HFB pairing field.
@@
The amount of pairing correlations present along the fission path from ground state to scission can be controlled in various ways. One may increase the absolute value of the pairing energy by raising the pairing gap ($\Delta$), the particle number fluctuations ($\Delta\hat{N}^{2}$), or the pairing strength.
@@
Conversely, the increase of the pairing strength by scaling the quenching factor leads to a decrease of the potential energy. Lowering the fission barrier and the collective inertias simultaneously leads to a monotonic reduction of the fission half-lives for stronger pairing interactions. The effect of lowering the barrier with an increase of quenching factor is less significant in the region of the asymmetric fission channel. Thus, enhancing the pairing strength has a different impact on fission properties compared to the two previous techniques. This is because the nuclear interactions are modified. Changes in $\Delta$ and $\Delta\hat{N}^{2}$ affects only the single-particle level properties around the Fermi level, but the nuclear force remains unchanged.
The effective inertias associated with the $\Delta\hat{N}^{2}$ computed within the “one-body assumption" are in very good agreement with the ones obtained for the pairing gap, which is a genuine one body operator. This result gives a solid ground to the use of $\Delta\hat{N}^{2}$ as collective variable in fission studies.